# Cardiocirculatory Computational Models for the Study of Hypertension

### v1

@@ 6 Conclusions
This work developed and applied a computational modeling framework to study hypertension in systemic, pulmonary and renovascular forms and in three nuances of severity, using both 0D models (with and without capillaries) [31] and a 3D–0D coupled model with electromechanical representation of the left ventricle [29], within a perspective that moved toward the development of cardiovascular DTs. The methodology included parameter calibration, global and multi-parametric sensitivity analysis and systematic comparison between models with and without capillaries.
@@
We recognize that our study has certain limitations. The calibration was conducted on synthetic data generated by $(\mathscr{C}_{\text{C}})$ model: it ensures a controlled comparison, but introduces the risk of not fully reflecting the individual variability of real patients. The lack of direct calibration of the 3D–0D model means that some mechanical aspects of the ventricle (particularly electrophysiological parameters and active contractility parameters) remain unoptimised and potentially less realistic, as certain physiological mechanisms regulating contraction may not be fully captured. 3D simulation requires high computational effort, limiting the number of scenarios that can be explored and its applicability in real-time clinical settings; this can be circumvented by means of reduced order models [65].
A significant extension of the work could arise from the integration of real clinical data to calibrate and validate the models. Using specific measurements, such as pressure, flow, volume, cardiac geometries, would allow the creation of true cardiovascular DTs, potentially useful for personalised diagnostics and monitoring. Such an approach requires more robust strategies to handle inter-individual variability and implement adaptive cost functions. Possible avenues of extension include 3D modeling of vascular and cardiac structures relevant in hypertension (e.g. [12, 30, 66]):
@@
    Renovascular hypertension with secondary pulmonary hypertension: add a 3D representation of the renal circulation, with dedicated arteries and veins, and the inclusion of equations describing the effect of sodium concentration on pressure control. The kidney contributes to hypertension both through diuretic and natriuretic mechanisms [67, 26] and by modulating sympathetic tone through reflexes that increase nerve activity and blood pressure [27]: these aspects are particularly important in salt-sensitive cases and in chronic conditions, particularly in renovascular hypertension.
@@ Appendix B Healthy individual parameters
Note: the parameters presented in Tables 59 are a modification of literature values [68, 69] in such a way that the model outputs lie in the healthy individual’s ranges, listed in C.
@@ Appendix C Healthy individual’s ranges
$\text{LA}_{\text{IVmax}}$ $[24,57]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{LA}_{\text{IVmin}}$ $[9,28]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{LA}_{\text{IVpreAC}}$ $[15,46]$ $\text{mL}/\text{m}^{2}$ [71]
$\text{LA}_{\text{PassEF}}$ $[8,44]$ % [71]
$\text{LA}_{\text{ActEF}}$ $[17,58]$ % [71]
$\text{LA}_{\text{TotEF}}$ $[37,70]$ % [70]
$\text{LA}_{\text{Pmax}}$ $[6,20]$ mmHg [72]
$\text{LA}_{\text{Pmin}}$ [2,9][-2,9] mmHg [72]
$\text{LA}_{\text{Pmean}}$ $[4,12]$ mmHg [72]
@@
$\text{RA}_{\text{IVmax}}$ $[28,76]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{RA}_{\text{IVmin}}$ $[9,45]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{RA}_{\text{IVpreAC}}$ $[19,61]$ $\text{mL}/\text{m}^{2}$ [71]
$\text{RA}_{\text{PassEF}}$ $[4,41]$ % [71]
$\text{RA}_{\text{ActEF}}$ $[11,55]$ % [71]
$\text{RA}_{\text{TotEF}}$ $[29,68]$ % [70]
$\text{RA}_{\text{Pmax}}$ $[2,14]$ mmHg [72]
$\text{RA}_{\text{Pmin}}$ [2,6][-2,6] mmHg [72]
$\text{RA}_{\text{Pmean}}$ [1,8][-1,8] mmHg [72]
@@
$\text{LV}_{\text{ISV}}$ $[30,66]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{LV}_{\text{IEDV}}$ $[47,107]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{LV}_{\text{IESV}}$ $[11,47]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{LV}_{\text{EF}}$ $[51,76]$ % [70]
$\text{LV}_{\text{Pmax}}$ $[90,140]$ mmHg [72]
$\text{LV}_{\text{Pmin}}$ $[4,12]$ mmHg [72]
@@
$\text{RV}_{\text{ISV}}$ $[28,75]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{RV}_{\text{IEDV}}$ $[53,123]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{RV}_{\text{IESV}}$ $[17,59]$ $\text{mL}/\text{m}^{2}$ [70]
$\text{RV}_{\text{EF}}$ $[42,72]$ % [70]
$\text{RV}_{\text{Pmax}}$ $[15,28]$ mmHg [72]
$\text{RV}_{\text{Pmin}}$ $[0,8]$ mmHg [72]
@@
CI $[2.8,4.2]$ $\text{m}^{2}\cdot\text{L}/\text{min}$ [72]
$\text{SAP}_{\text{max}}$ $[0,140]$ mmHg [73]
$\text{SAP}_{\text{min}}$ $[0,80]$ mmHg [73]
$\text{PAP}_{\text{max}}$ $[15,28]$ mmHg [72]
$\text{PAP}_{\text{min}}$ $[5,16]$ mmHg [72]
$\text{PAP}_{\text{mean}}$ $[10,22]$ mmHg [72]
$\text{PWP}_{\text{max}}$ $[9,23]$ mmHg [72]
$\text{PWP}_{\text{min}}$ $[1,12]$ mmHg [72]
$\text{PWP}_{\text{mean}}$ $[6,15]$ mmHg [72]
SVR $[11.3,17.5]$ $\text{mmHg}\cdot\text{min}/\text{L}$ [72]
PVR $[1.9,3.1]$ $\text{mmHg}\cdot\text{min}/\text{L}$ [72]
$\text{S}_{f}$ $[0,0.05]$ $1/\text{m}^{2}$ [74]
@@
Note: the ranges presented in Tables 1114 are derived from echocardiography (the ones taken from [73, 72]) and magnetic resonance imaging (MRI) (the ones taken from [70, 71]), meaning that using general healthy ranges may not always be entirely appropriate; however, this is not an issue since, despite the differences between echocardiography and MRI, the quantities computed by both $(\mathscr{C}_{\text{NC}})$ and $(\mathscr{C}_{\text{C}})$ for a healthy individual fall within all ranges.

### v2

@@ 6 Conclusions
This work developed and applied a computational modeling framework to study hypertension in systemic, pulmonary and renovascular forms and in three nuances of severity, using both 0D models (with and without capillaries) [lumped:capillary] and a 3D–0D coupled model with electromechanical representation of the left ventricle [3d-0d_ventr], within a perspective that moved toward the development of cardiovascular DTs. The methodology included parameter calibration, global and multi-parametric sensitivity analysis and systematic comparison between models with and without capillaries.
@@
We recognize that our study has certain limitations. The calibration was conducted on synthetic data generated by $(\mathscr{C}_{\text{C}})$ model: it ensures a controlled comparison, but introduces the risk of not fully reflecting the individual variability of real patients. The lack of direct calibration of the 3D–0D model means that some mechanical aspects of the ventricle (particularly electrophysiological parameters and active contractility parameters) remain unoptimised and potentially less realistic, as certain physiological mechanisms regulating contraction may not be fully captured. 3D simulation requires high computational effort, limiting the number of scenarios that can be explored and its applicability in real-time clinical settings; this can be circumvented by means of reduced order models [ldnets].
A significant extension of the work could arise from the integration of real clinical data to calibrate and validate the models. Using specific measurements, such as pressure, flow, volume, cardiac geometries, would allow the creation of true cardiovascular DTs, potentially useful for personalised diagnostics and monitoring. Such an approach requires more robust strategies to handle inter-individual variability and implement adaptive cost functions. Possible avenues of extension include 3D modeling of vascular and cardiac structures relevant in hypertension (e.g. [model_hyper:es, model-whole-heart1, model-two-chambers]):
@@
    Renovascular hypertension with secondary pulmonary hypertension: add a 3D representation of the renal circulation, with dedicated arteries and veins, and the inclusion of equations describing the effect of sodium concentration on pressure control. The kidney contributes to hypertension both through diuretic and natriuretic mechanisms [kidney7, kidney6] and by modulating sympathetic tone through reflexes that increase nerve activity and blood pressure [kidney8]: these aspects are particularly important in salt-sensitive cases and in chronic conditions, particularly in renovascular hypertension.
@@ Acknowledgements
The authors acknowledge Dr. Andrea Parlangeli for having facilitated the collaboration between Professors Parati and Quarteroni on this interesting scientific topic.
@@ Appendix B Healthy individual parameters
Note: the parameters presented in Tables 59 are a modification of literature values [healthy-params-fonte1, healthy-params-fonte2] in such a way that the model outputs lie in the healthy individual’s ranges, listed in C.
@@ Appendix C Healthy individual’s ranges
$\text{LA}_{\text{IVmax}}$ $[24,57]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{LA}_{\text{IVmin}}$ $[9,28]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{LA}_{\text{IVpreAC}}$ $[15,46]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges4]
$\text{LA}_{\text{PassEF}}$ $[8,44]$ % [healthy-ranges4]
$\text{LA}_{\text{ActEF}}$ $[17,58]$ % [healthy-ranges4]
$\text{LA}_{\text{TotEF}}$ $[37,70]$ % [healthy-ranges3]
$\text{LA}_{\text{Pmax}}$ $[6,20]$ mmHg [healthy-ranges2]
$\text{LA}_{\text{Pmin}}$ [2,9][-2,9] mmHg [healthy-ranges2]
$\text{LA}_{\text{Pmean}}$ $[4,12]$ mmHg [healthy-ranges2]
@@
$\text{RA}_{\text{IVmax}}$ $[28,76]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{RA}_{\text{IVmin}}$ $[9,45]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{RA}_{\text{IVpreAC}}$ $[19,61]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges4]
$\text{RA}_{\text{PassEF}}$ $[4,41]$ % [healthy-ranges4]
$\text{RA}_{\text{ActEF}}$ $[11,55]$ % [healthy-ranges4]
$\text{RA}_{\text{TotEF}}$ $[29,68]$ % [healthy-ranges3]
$\text{RA}_{\text{Pmax}}$ $[2,14]$ mmHg [healthy-ranges2]
$\text{RA}_{\text{Pmin}}$ [2,6][-2,6] mmHg [healthy-ranges2]
$\text{RA}_{\text{Pmean}}$ [1,8][-1,8] mmHg [healthy-ranges2]
@@
$\text{LV}_{\text{ISV}}$ $[30,66]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{LV}_{\text{IEDV}}$ $[47,107]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{LV}_{\text{IESV}}$ $[11,47]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{LV}_{\text{EF}}$ $[51,76]$ % [healthy-ranges3]
$\text{LV}_{\text{Pmax}}$ $[90,140]$ mmHg [healthy-ranges2]
$\text{LV}_{\text{Pmin}}$ $[4,12]$ mmHg [healthy-ranges2]
@@
$\text{RV}_{\text{ISV}}$ $[28,75]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{RV}_{\text{IEDV}}$ $[53,123]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{RV}_{\text{IESV}}$ $[17,59]$ $\text{mL}/\text{m}^{2}$ [healthy-ranges3]
$\text{RV}_{\text{EF}}$ $[42,72]$ % [healthy-ranges3]
$\text{RV}_{\text{Pmax}}$ $[15,28]$ mmHg [healthy-ranges2]
$\text{RV}_{\text{Pmin}}$ $[0,8]$ mmHg [healthy-ranges2]
@@
CI $[2.8,4.2]$ $\text{m}^{2}\cdot\text{L}/\text{min}$ [healthy-ranges2]
$\text{SAP}_{\text{max}}$ $[0,140]$ mmHg [healthy-ranges1]
$\text{SAP}_{\text{min}}$ $[0,80]$ mmHg [healthy-ranges1]
$\text{PAP}_{\text{max}}$ $[15,28]$ mmHg [healthy-ranges2]
$\text{PAP}_{\text{min}}$ $[5,16]$ mmHg [healthy-ranges2]
$\text{PAP}_{\text{mean}}$ $[10,22]$ mmHg [healthy-ranges2]
$\text{PWP}_{\text{max}}$ $[9,23]$ mmHg [healthy-ranges2]
$\text{PWP}_{\text{min}}$ $[1,12]$ mmHg [healthy-ranges2]
$\text{PWP}_{\text{mean}}$ $[6,15]$ mmHg [healthy-ranges2]
SVR $[11.3,17.5]$ $\text{mmHg}\cdot\text{min}/\text{L}$ [healthy-ranges2]
PVR $[1.9,3.1]$ $\text{mmHg}\cdot\text{min}/\text{L}$ [healthy-ranges2]
$\text{S}_{f}$ $[0,0.05]$ $1/\text{m}^{2}$ [shunt-ranges]
@@
Note: the ranges presented in Tables 1114 are derived from echocardiography (the ones taken from [healthy-ranges1, healthy-ranges2]) and magnetic resonance imaging (MRI) (the ones taken from [healthy-ranges3, healthy-ranges4]), meaning that using general healthy ranges may not always be entirely appropriate; however, this is not an issue since, despite the differences between echocardiography and MRI, the quantities computed by both $(\mathscr{C}_{\text{NC}})$ and $(\mathscr{C}_{\text{C}})$ for a healthy individual fall within all ranges.